OpenStudy (anonymous):

Help please! A rectangle has area 16m^2. Express the perimeter of the rectangle as a function of the length of one of its sides.

OpenStudy (anonymous):

\[ A =l\times w = 16 \implies w = \frac{16} l \]\[ P = 2l + 2w = 2l + 2\frac{16} l \]

OpenStudy (anonymous):

Simplify.